Beam Delivery Report

From exp-prog

Jump to: navigation, search

Status reports on beam delivery at TRIUMF, and part of the Beam Delivery group's Guide for Experimenters.

Past reports can be found here.

Permalinks: Cyclotron status Cyclotron log ISAC log

[edit] 18:45, 24 August 2010 (PDT)

SERVICE UPDATE: Because the wiki framework doesn't lend itself well to blog-like periodic entries, this page is being deprecated. In its place will be a Beam Delivery Updates weblog on TRIUMF's documents server. Past entries will be maintained here, but future entries will be posted there on an as-needed basis. The main page will be updated to direct readers to the new location. — Morton

[edit] 17:41, 23 August 2010 (PDT)

The cyclotron is running well. After a brief interruption this morning to investigate a leak in the copper active water system, the cyclotron has continued to deliver ~125 µA to BL1A, 45 µA to BL2A, and ~75 µA to BL2C4. There are a number of interruptions to different beamlines planned over the next few days. There will be maintenance at ISAC tomorrow (BL1A will remain online), a brief period of cyclotron development Wednesday (as high-current operation with BL1A ends for Schedule 117), a change of the 1A extraction foil cartridge Thursday in preparation for low-current operation with BL1A/B, and patient treatment over the weekend.

At ISAC, TUDA experiment S1233 has been cancelled due to a failure of the Phase 1 (cryogenics) compressor. We'll try to start the b-NMR experiments originally scheduled for Friday a few shifts early and use a shift in the interim for TRILIS development. The online schedule has been updated to reflect these changes. — Morton

[edit] 22:07, 13 August 2010 (PDT)—WEEKEND UPDATE

Everything is running relatively well. The cyclotron is currently delivering ~125 µA to BL1A, 45 µA to BL2A, and ~77 µA to BL2C4. This is expected to continue through the weekend.

At ISAC, 30Na is being delivered to the Osaka experiment and data-taking is underway. Polarization was successfully established earlier in the week. That experiment will continue until next Wednesday's scheduled maintenance day. At ISAC-II, SCC commissioning is underway using stable beam from OLIS. That will also continue next week. — Morton

[edit] 22:42, 10 August 2010 (PDT)

After a relatively uneventful weekend, the cyclotron is delivering ~125 µA to BL1A, 45 µA to BL2A, and ~75 µA to BL2C4. This will continue overnight. Maintenance is scheduled for tomorrow; beam is expected to be unavailable from 09:00 until 13:00 or so. Please note that the cyclotron will remain on for the first couple of hours as the current machine tune is documented; this is follow-up work from last week's development shift.

30Na has been delivered to the Osaka experiment at ISAC, but has proven difficult to polarize. They're now taking 26Na to try to establish polarization before switching back. That experiment will continue into next week. In addition, OLIS is being used in support of SCC commissioning. This will also continue through the week. — Morton

[edit] 17:15, 4 August 2010 (PDT)

Beam development will continue into the night, with maintenance (including an ISIS backplate change) planned for tomorrow. The full maintenance list is here. At this time, beam is expected late Thursday evening.

Stable-beam tuning from the new low-power Ta target in ITW is underway. The high-power ZrC target used for development over the past four weeks has reached end-of-life and will be disconnected and removed over the next two weeks. Protons are expected on the new target following maintenance, with delivery to experiment (30Na for Osaka experiment S1114) planned for this weekend. — Morton

[edit] 17:37, 30 July 2010 (PDT)—(LONG) WEEKEND UPDATE

BL1A is getting a little bit more beam than yesterday (~130 µA) while BL2A remains at 75 µA. Beam is still being delivered to BL2C1 for PIF (at the nA level); this is scheduled to continue overnight with isotope production on BL2C4 resuming afterwards. We expect to run ~125/75/75 µA to BL1A/2A/2C4 over the rest of the weekend.

Yield measurements and source development are continuing at ISAC. Protons will be required off for several hours tomorrow (Saturday) to accommodate the move of TM1 to ITW; this was originally planned for last Thursday, but a leak in a cooling line delayed it. Development work will continue over the weekend, with an attempt to accelerate a charge-bred beam to SEBT3A/TIGRESS planned for Tuesday. — Morton

[edit] 18:06, 29 July 2010 (PDT)

The cyclotron is currently delivering ~125 µA to BL1A, 75 µA to BL2A, and tens of nA (off and on) to BL2C1 for PIF. This should continue overnight.

At ISAC, target/source/CSB development is continuing. A decision as to whether or not we'll try to accelerate a charge-bred beam over the weekend will be made tomorrow. — Morton

[edit] 17:47, 27 July 2010 (PDT)

Beam was restored overnight, then stopped again due to RF problems, then restored again this morning, before BL2A had to be taken offline due to problems with the 2AQ15 power supply. That all seems to be sorted out now and the cyclotron is delivering ~120 µA to BL1A, 75 µA to BL2A, and tens of nA to BL2C1. This should continue overnight, with periodic accesses to 2C1 for PIF.

At ISAC, 4He is being tuned to TIGRESS for calibration purposes. This is being held up due to a problem getting the Prague magnet (used to determine the beam energy from the DTL) power supply to come on. Experts are being contacted. On the RIB side, CSB development and yield measurements are underway in preparation for a production test and delivery to TIGRESS later this week. — Morton

[edit] 18:55, 26 July 2010 (PDT)

Back from vacation. Things were running well until about 16:00, at which point a power bump knocked a bunch of things offline. Injection is expected around midnight, with proton delivery some time after that. The plan is to deliver ~125 µA to BL1A, 75 µA to BL2A, and ~0.1 µA to BL2C1 for PIF. We'll see how things look tomorrow morning. — Morton

[edit] 23:06, 30 June 2010 (PDT)

The cyclotron is currently delivering beam to BL1A and 2C4 (~115 µA and ~75 µA, respectively), but BL2A is off for the foreseeable future. Efforts to deliver RIB from the current target have been abandoned and investigations into the HV problems that have plagued us are underway. The scheduled experimenters will try to run with stable beam from OLIS instead.

This will be the last update for a while as I will be on vacation until July 26. As always, check the cyclotron and ISAC logs and the cyclotron status page (see above) for up-to-date information. — Morton

[edit] 22:58, 26 June 2010 (PDT)—WEEKEND UPDATE

The current has been reduced in BL2A, but the cyclotron is still delivering beam to all three high-current beamlines: ~115 µA to BL1A, ~75 µA to BL2C4, and 40 µA to BL2A. We're not sure about BL2A, but the other currents are expected to remain at roughly those values until Monday.

We've had good news and bad news at ISAC. The good news? Despite our misgivings, we were able to reach 49 kV with the target bias overnight Friday, allowing us to start work towards delivering 24Na without the CSB. The bad news? It didn't hold—the target began sparking early afternoon, and we've been trying to restore the bias voltage ever since. Ops will continue to do so overnight, but we may end up having to attempt delivery with the CSB after all. Again, we'll know more in the morning. — Morton

[edit] 18:48, 25 June 2010 (PDT)—WEEKEND UPDATE

We're finally running beam to all three beamlines: ~105 µA to BL1A, ~75 µA to BL2C4, and 60 µA to BL2A. We expect to continue like this overnight, but may raise the current in BL2A to 70 µA over the weekend.

We ended up abandoning the attempt to run ISAC at high bias voltages due to sparking. This cost us experiment S1204 and led us to attempt delivery of 24Na to TIGRESS using the Charge State Booster. (With the CSB we can get to 24Na5+, allowing us to achieve the required 2.04 keV/u needed for injection into the RFQ at a bias voltage of 9.79 kV.) Unfortunately this resulted in a 24Na rate well below that required for the experiment. We're taking one last stab at running at high bias voltage now that the target has seen protons and had a few days to outgas, but if that doesn't work we'll attempt delivery with the CSB (as a development exercise, as much as anything). We'll know more in the morning. — Morton

[edit] 20:29, 21 June 2010 (PDT)

We're currently delivering ~120 µA to BL1A and ~75 µA to BL2C4. The 2C4 target was changed earlier today to meet a delivery deadline. There has been some progress in understanding last week's TNF protect monitor beam trips—a correlation has been identified with the 1AVQ2 current—so now it's a matter of identifying and implementing a solution. We'll continue to run at these currents overnight. Tomorrow is a maintenance day (scheduled for 24 hours, but expected to be ~10) to allow the replacement of the ISIS backplate.

At ISAC, things are still going slowly. The shielding blocks over the west target station were removed to allow access to the top of the target module and evidence of sparking was found under the HV cover. Repairs have been made and Ops is attempting, once again, to raise the target bias voltage to above 50 kV. Protons will not be put on target overnight. A decision regarding the beam schedule beyond tomorrow's scheduled maintenance still has to be made. — Morton

[edit] 22:57, 19 June 2010 (PDT)

The cyclotron is running relatively well, but ISAC remains offline as the target will not yet hold the 53.04 kV needed for 26Al delivery. Conditioning will continue overnight. — Morton

[edit] 18:18, 18 June 2010 (PDT)—WEEKEND UPDATE

The cyclotron is currently delivering ~120 µA to BL1A and a few nA to BL2C1 for PIF. Isotope production on BL2C4 (at ~75 µA or so) will resume once PIF is finished this evening. BL2A will remain offline overnight.

The target leak at ISAC has been repaired and HV and target heaters are being restored. Assuming all goes well overnight, the shielding over the target station will be put back in place tomorrow morning and protons put on target. It will take some time for the proton current to be ramped up to 70 µA. Yield measurements and TRILIS development will likely begin long before that, with the hope of delivering some beam to experiment (even if we're still ramping up) over the weekend. — Morton

[edit] 16:15, 17 June 2010 (PDT)

ISAC remains offline as a vacuum leak in the west target station is investigated. Beam will not be restored to BL2A before tomorrow. In the meantime the cyclotron is delivering ~120 µA to BL1A and a few nA to BL2C1. This is scheduled to continue until Saturday morning. — Morton

[edit] 17:59, 16 June 2010 (PDT)

After suffering frequent TNF protect monitor trips over the weekend, the cyclotron has been relatively stable since Monday. It's currently delivering ~115 µA to BL1A and a few nA (off and on) to BL2C1 for PIF. After taking up to 25 µA over the last few days, ISAC is offline due to an apparent vacuum leak. That will be investigated further tomorrow after the fields in the module access area have decayed. The scheduled ISAC experiment will take stable beam from OLIS in the interim. — Morton

[edit] 15:31, 11 June 2010 (PDT)—WEEKEND UPDATE

The cyclotron is currently delivering ~120 µA to BL1A and ~75 µA to BL2C4. There have been a number of beam trips due to instability in the position of the 1A beam at TNF; Ops and beam experts have been working to identify the source of the problem for most of the morning.

ISAC is still offline. The new production target is being conditioned; the hope is to get some beam (~10 µA) on target Sunday so that target scans and beam development can begin Tuesday. TIGRESS will continue to take beam from OLIS until Saturday morning at which point it, too, will be taken offline in preparation for the next experiment. — Morton

[edit] 15:36, 8 June 2010 (PDT)

The cyclotron is running and is relatively stable, delivering ~120 µA to BL1A and ~70 µA to BL2C4. BL2A is off until the next ISAC target is ready for beam (likely this weekend). After a brief interruption this morning to access STF, during which the B20 tripped off (and was restored), we expect no interruptions over the next few days. There will be no maintenance period tomorrow.

At ISAC, 15N is being set up to TIGRESS. They'll take beam until Saturday morning at which point we'll switch from the Supernanogan back to the microwave source for the next few experiments. — Morton

[edit] 20:41, 7 June 2010 (PDT)

The RF is being brought up. It's currently at 90 kV; once it's at 95 kV and stable, injection and cyclotron tuning will commence. Beam delivery may be possible some time overnight. Keep checking the cyclotron status page and log for more information. — Morton

[edit] 14:56, 7 June 2010 (PDT)

The B20 has been brought online more quickly than anticipated. There is a chance beam might be restored late this evening or early tomorrow morning. — Morton

[edit] 09:55, 7 June 2010 (PDT)

The Linde has still not been returned to operation. It's possible that the cryogenics group will switch back to the old B20 refrigerator to allow the cyclotron to run while the problem is addressed. That decision has not yet been made. More information should be available at this morning's 11:00 Ops/performance meeting (in the Auditorium).

At ISAC a beam of 15N4+ from the Supernanogan offline source has been developed. Accelerator tuning (to TIGRESS) is planned for today. — Morton

[edit] 23:26, 4 June 2010 (PDT)—WEEKEND UPDATE

The Linde has been shut down. The thinking is that there is still contamination in the system that can only be addressed at room temperature. Work to restore the refrigerator will resume tomorrow morning. There is no estimate as to when beam will be restored. — Morton

[edit] 16:50, 4 June 2010 (PDT)

The cryogenics experts are still working to get the Linde refrigerator running stably. Until they succeed, we have no idea when beam will be restored. Cyclotron Ops will continue to post updates here and here.

At ISAC, RIB delivery to β-NMR has been extended until Sunday morning. Stable-beam delivery to DRAGON is still scheduled to end tomorrow morning. — Morton

[edit] 13:51, 4 June 2010 (PDT)

The Linde 1630 helium refrigerator was restarted this morning and the cryopanels are cooling. We still don't have an estimate as to when beam might be injected into the cyclotron. — Morton

[edit] 17:38, 3 June 2010 (PDT)

The Linde began to act up coming out of yesterday's maintenance. It was warmed up this morning and will be restarted tomorrow (Friday). No beam is expected before 18:00 Friday. For more information, check the cyclotron status page or the cyclotron and ISAC logs. — Morton

[edit] 14:49, 2 June 2010 (PDT)

Maintenance is underway. Protons are expected to be restored early this evening, with delivery to experiments to follow.

The ISAC beam schedule has been updated. See here for more information. TM3 has been repaired and is currently being tested in the ISAC hot cell; it's expected to be moved back to ITW tomorrow. This will take 5–6 hours during which RIB will not be available. The exact time has yet to be determined. — Morton

[edit] 14:36, 1 June 2010 (PDT)

Cyclotron development is underway. The plan is to develop a tune with BL2A operating at 480 MeV and BL1A operating at 500 MeV and to run like that overnight. The previous 480/480 MeV tune will be restored prior to tomorrow's maintenance.

An ISIS backplate change is planned for tomorrow's maintenance. While maintenance is scheduled for 24 hours, beam is expected overnight Wednesday (barring last-minute demands for work).

The ISAC schedule is about to be changed as the extraction electrode on the next target failed during high-voltage conditioning, forcing a trip back to the hot cell for repairs. The current experiments will continue until Saturday morning. — Morton

[edit] 13:57, 25 May 2010 (PDT)

It's been an unusual day in that BL2A is off for ISAC maintenance and BL2C4 was off for a target change while BL1A continued to take beam uninterrupted. We're currently delivering ~125 µA to BL1A and ~70 µA to BL2C4; we'll resume delivering 40 µA to BL2A later this afternoon. We expect to run like this through the week.

At ISAC, β-NMR will resume data-taking with 8Li once protons are restored. The off-line source is being set up to deliver 33S from the Supernanogan to DRAGON later this week. — Morton

[edit] 15:46, 18 May 2010 (PDT)

Maintenance is well underway. The cyclotron RF is back on and beam is expected to be restored some time this evening. At ISAC, stable beam tuning (from the offline ion source) to TIGRESS is underway; online tuning to βNMR will begin once beam has been restored. — Morton

[edit] 16:12, 17 May 2010 (PDT)

The cyclotron's currently delivering ~100 µA to BL1A, 70 µA to BL2A, and ~70 µA to BL2C. This should continue until tomorrow's maintenance. That will take most of the day as an ISIS backplate change is planned; however, proton delivery is expected overnight.

At ISAC, S1147 will wrap up with the start of tomorrow's maintenance. TRILIS development and setup for the upcoming β-NMR experiment(s) will follow, likely at a reduced proton current. The full proton current will be required once we deliver to the experiment on Thursday. A 6–8 hour interruption will be required on Friday to accommodate a target module move. — Morton

[edit] 21:55, 14 May 2010 (PDT)—WEEKEND UPDATE

The T1 water leak has been repaired (and there are photos in the cyclotron log) and beam is being restored to BL1A. The currents on BL2A and BL2C4 will be reduced while 1A is brought back online. In the meantime we've been delivering 65 µA to BL2A and ~70 µA to BL2C4; the hope is to restore those and bring 1A back up to ~110 µA as quickly as possible. Otherwise, no changes are expected through the weekend. — Morton

[edit] 13:36, 13 May 2010 (PDT)

BL1A is offline due to a water leak at the T1 target. Shielding blocks are currently being removed to access T1. The full course of action won't be known until the source of the leak is identified. In addition, the beam profile monitor at T2 has failed; that monitor would be needed to deliver beam to T2, TNF, and the 500 MeV irradiation facility with T1 removed for repairs. Once the blocks have been removed from above T1 work will begin to excavate T2.

In the meantime, the cyclotron continues to deliver 65 µA to BL2A and ~70 µA to BL2C4. The ISAC schedule has been updated again: S1147 will run until 08:30 Tuesday and S1142 (146Tb at the 8π) has been cancelled. β-NMR will run an extra week in its place. — Morton

[edit] 21:08, 12 May 2010 (PDT)

The cyclotron is currently delivering ~105 µA to BL1A, 65 µA to BL2A, and ~70 µA to BL2C4. We lost 42 minutes to an RF crowbar this morning, but otherwise things have been relatively stable. We expect to continue running like this through the weekend.

11Li delivery to experiment S1147 continues at ISAC. The schedule has been changed to allow them to run without interruption until Monday. — Morton

[edit] 14:37, 11 May 2010 (PDT)

We lost time yesterday afternoon due to an unrelated problem with the Linde helium refrigerator—a fuse blew, causing it to trip off. That was restored yesterday evening and we've been running relatively smoothly since. The RF is still sparking occasionally, but the machine is gradually settling down.

We're currently delivering ~110 µA to BL1A, 65 µA to BL2A, and ~65 µA to BL2C4. We'd like the currents in BL1A and BL2C4 to be a bit higher, but the machine transmission is still not great (though it's much better than it was yesterday or over the weekend). Tomorrow's maintenance has been cancelled outright (see the beam schedule and maintenance list), so we expect to continue delivering high currents until next week. The RIB experiment currently running at ISAC will continue until Monday morning as well, with TRILIS development to follow. — Morton

[edit] 10:49, 10 May 2010 (PDT)

After all of last week's problems we were able to run relatively smoothly over the weekend. The cyclotron is currently delivering ~110 µA to BL1A, 65 µA to BL2A, and ~65 µA to BL2C4. Transmission through the machine is still poor, though, so some tuning work is likely required.

The current state of the machine and plans for the coming week will be discussed in some detail at the weekly Ops/performance meeting (11:00 in the Auditorium). All are encouraged to attend. — Morton

[edit] 22:59, 7 May 2010 (PDT)—WEEKEND UPDATE

The dee voltage is at 90 kV, the Linde motor frequency is still relatively stable—there's certainly no evidence of the fast oscillations we were seeing all week—and beam has been injected. There are currently 30+ µA being extracted down BL1A, with BL2A and 2C4 to follow. The transmission is still poor (not surprising, given that the machine has been off for most of the last six days) and the machine will likely need several hours to settle down, but it's a start.

Should the cryogenic system hold, the plan is to run high currents in all three beamlines over the weekend: BL1A is always run at >100 µA, while ISAC will want 65 µA in BL2A and BL2C4 will need 60–70 µA for isotope production. It will likely take some time to bring the total extracted current to the 245 µA or so that we'd like. In addition, there may be periods of reduced beam intensity to accommodate cyclotron tuning. We appreciate how patient everyone has been through this, but we still need a few more days to get back to where we should be.

Thanks. — Morton

[edit] 18:07, 7 May 2010 (PDT)

And here we go again: The RF dee voltage is at 70 kV and (so far) the Linde appears to be holding steady. Ops has been asked not to increase the dee voltage faster than ~10 kV/hour, so it would still be ~20:00 or so before they'd be at 90 kV and some time after that before they'd be in a position to try injecting into the cyclotron. The plan is to deliver beam if possible, though the current may be limited as the RF settles down. — Morton

[edit] 14:34, 7 May 2010 (PDT)

Things are still a mess—last night's attempt to restore the RF failed as the Linde motor frequency became unstable (again). An engineer from Linde is en route from Tulsa, Oklahoma; he'll be on site tomorrow morning to try to diagnose the problem. In the meantime, we're going to keep trying to get things going. Another attempt is being made to bring the RF online in the hopes that the cryogenics will remain stable, but even if this attempt is successful there is clearly a problem which will still have to be addressed. — Morton

[edit] 19:42, 6 May 2010 (PDT)

After defrosting overnight, the cyclotron cryogenics system was restarted this morning. The tank vacuum is good and the system appears to be stable. The RF was started about half an hour ago and the dee voltage is slowly being brought up to its operating value of 90 kV. If things remain stable, injection and extraction down BL1A will follow, with extraction down BL2C4 and 2A after that. Again, check the cyclotron and ISAC logs for more information. — Morton

[edit] 15:03, 5 May 2010 (PDT)

Beam will be off until at least tomorrow morning as the cryopanel system is brought back to room temperature and then re-cooled. We're taking the opportunity to perform unrelated maintenance, including the ISIS backplate change that was going to be necessary next week. Hopefully we'll know more by tomorrow. — Morton

[edit] 08:34, 5 May 2010 (PDT)

And we're down again. We had beam briefly overnight, but the same problems resurfaced around 3:00 this morning. As usual, we have no idea how long it will take to resume beam—until the cryogenics and RF groups give the okay to restart the RF, we can't make any estimates

If you have any maintenance jobs that might take a few hours but can be wrapped up in an hour if the word is given to restart the RF, today would probably be a good day to do them. Contact Ops if you need a work permit; they'll be able to advise. — Morton

[edit] 15:24, 4 May 2010 (PDT)

Work on the Linde is continuing, but there's still no reliable estimate of how long it will be until the problem is under control. Ops is posting updates on the Cyclotron Status page; additional information may be found in the Cyclotron and ISAC logs. — Morton

[edit] 08:14, 4 May 2010 (PDT)

The good news? We were able to deliver beam overnight. The bad news? The same issues we've been having with the Linde resurfaced overnight. The RF has been off since ~07:00 this morning. We don't know how long it will be until it's restored.

The vacuum and RF experts are working on the problem in consultation with outside experts. Once the cryogenic system is stable, we expect a startup like yesterday's (i.e. a few hours from injection to full current in all beamlines). — Morton

[edit] 17:57, 3 May 2010 (PDT)

As you may have noticed, the cyclotron is operating and beam has been restored. We're currently delivering ~115 µA down BL1A and 65 µA down BL2A, though this may be reduced for periods overnight and through tomorrow to allow Ops to work on the cyclotron tune. The Linde unit that supplies cold He gas to the cryopanels appears to be holding.

RIB tuning has resumed at ISAC. The hope is to switch from tuning to delivery of 11Li to experiment S1147 within the next few hours. — Morton

[edit] 12:45, 3 May 2010 (PDT)

While we were able to restore beam for a couple of days following last week's repair of the extraction probe, we have been unable to deliver beam since Saturday night due to problems with the cyclotron cryogenics system. (Large cryogenically cooled panels are used to trap hydrogen and other trace gases as part of the cyclotron vacuum system; without reliable cryogenics, we don't have reliable vacuum.) A final effort is being made, in consultation with outside experts, to bring the system into regulation in the hopes of restoring cyclotron RF this afternoon (with injection to follow some time after that), but at this point we still don't have a good estimate of how long it will be before beam delivery resumes. — Morton

[edit] 20:55, 29 April 2010 (PDT)

The cyclotron is still offline following yesterday's failure of the BL1A extraction probe. During a routine foil change the old extraction foil hung up on the probe mechanism; the cyclotron had to be vented for that foil to be removed before the probe could be withdrawn to its housing and repaired. The tank is back under vacuum and the probe is currently being tested. If all goes well, we expect proton beam delivery to resume tomorrow.

At ISAC, S1151 wraps up tonight. We begin stable beam tuning for the upcoming RIB experiments at ISAC-II tomorrow. We also hope to bring the new production target online, with target scans tomorrow and yield measurements over the weekend. RIB delivery to experiment S1147 is scheduled for Monday. — Morton

[edit] 13:40, 27 April 2010 (PDT)

The cyclotron is currently delivering ~125 µA to BL1A and ~75 µA to BL2C4. This is expected to continue overnight, with maintenance, beam development, and training planned for tomorrow. Regular beam delivery is scheduled to resume 08:30 Thursday with BL2A also coming online.

16O delivery to experiment S1151 is continuing at ISAC-II. That run will end Friday morning with stable beam tuning for two upcoming RIB experiments at ISAC-II to follow. — Morton

[edit] 10:48, 23 April 2010 (PDT)—WEEKEND UPDATE

The second round of patient treatments is complete. PT will continue Saturday and Sunday at 08:30. There are no interruptions planned for Monday; usually PT starts on a Friday (with setup on Thursday) so that the standard four rounds of treatment finishes the following Monday, but this week's treatment started a day earlier than usual. In the meantime, the cyclotron continues to deliver ~125 µA to BL1A. BL2C4 will come back online Sunday, with BL2A expected to remain offline until after Wednesday's maintenance.

Stable beam delivery is still planned at ISAC, with experiment S1151 scheduled to start Sunday. — Morton

[edit] 22:33, 20 April 2010 (PDT)

Cyclotron development was held up this morning when it became necessary to replace the ISIS backplate on short notice. That had originally been planned for next week's 24-hour maintenance. Beam is expected to be restored to BL1A some time tonight. BL2A will remain offline until next week.

Patient treatment is scheduled on BL2C1 for the next five days (though only setup is planned for Wednesday). There are two patients, so expect ~2 hours each morning in which beam will not be available in BL1A. — Morton

[edit] 16:46, 19 April 2010 (PDT)

The cyclotron is currently delivering ~115 µA to BL1A and ~75 µA to the Solid Target Facility (STF) on BL2C4. ISAC will remain offline (mostly—see below) until next week.

The beam development shift scheduled for April 28 has been moved to tomorrow (April 20) so that high-current development can be carried out using all three beamlines. The west target station at ISAC will be kept available until that is finished, then shut down so that the spent target can be removed next week. The 24-hour maintenance scheduled for April 29 will be moved to April 28 so that startup of the next ISAC production target can start a day earlier than originally planned; this will also prevent us having to start the cyclotron back up (following the scheduled ion source filament change) on a Friday. Otherwise the schedule will remain as-is, with the next experiment to take RIB at ISAC being experiment S1147 at TUDA-II.

The up-to-date beam schedule can be found here. — Morton

[edit] 23:04, 18 April 2010 (PDT)

ISAC is currently off-line as the ionizer on the RIB production target failed yesterday morning. There will be meetings tomorrow (Monday) to determine a course of action and what, if any, changes will be made to the schedule. The new target is not scheduled to come on-line until April 30. In the meantime, the cyclotron continues to deliver ~120 µA to BL1A and ~70 µA to BL2C4. This is not likely to change before Tuesday's "maintenance if required" shift. — Morton

[edit] 21:12, 16 April 2010 (PDT)—WEEKEND UPDATE

The cyclotron has been a bit unstable since about 13:00 today, but we're continuing to deliver ~110/70/70 µA to BL1A/2A/2C4, respectively. We'd still like about another 10 µA from the machine, but otherwise there are no plans to change things over the weekend.

S1007 continues to run at ISAC. They plan to take beam over the weekend as they've been doing, with long implant/decay counting cycles controlled manually by ISAC Ops. They may switch to shorter cycles on Monday, but otherwise we expect no changes before next Tuesday's if-required maintenance. — Morton

[edit] 17:29, 15 April 2010 (PDT)

Things are running much more smoothly than they were earlier in the week. We're currently delivering ~110/70/70 µA to BL1A/2A/2C4, respectively. Our goal is to maintain a total extracted current of 260 µA, so we're close (and have exceeded that off-and-on all day).

At ISAC, 115Ag is being delivered to experiment S1007 at the 8π. This will continue until next Tuesday. On the stable beam side, ISAC-II development and commissioning continues. — Morton

[edit] 18:53, 9 April 2010 (PDT) (WEEKEND UPDATE)

Beam was restored this morning, but the cyclotron is still not happy—the transmission is poor, and that's limiting the amount of beam that can be extracted. We've been delivering ~100/60/60 µA to BL1A/2A/2C4 all day, about 15% less than we'd like. Now that the machine's had a chance to settle down—the RF hasn't yet been on for 24 hours since recovering from yesterday's power bump—efforts will be made to tune the cyclotron this evening in the hopes of increasing the current in all three beamlines.

Both RIB and stable beam are being delivered at ISAC, 8Li to M1226 (βNQR) and 16O to DRAGON for a study of the acceptance of the spectrometer. Both of these are scheduled to continue over the weekend. On the RIB side, TRILIS development and setup to the 8π will follow; on the stable beam side, ISAC-II development. — Morton

[edit] 17:48, 8 April 2010 (PDT)

We're still waiting on cyclotron vacuum and RF. Otherwise things are largely recovered—beamline vacuums are good, ISIS is running (with beam extracted to the fast target), ISAC is ready to go, and experts are working on the stable beam tune through the ISAC accelerators for tomorrow's DRAGON test run. The estimate earlier today was for beam in BL1A around 20:00 or 21:00 this evening (with BL2A to follow); that may be optimistic, but things are coming together. Again, check the logs for more information. — Morton

[edit] 09:03, 8 April 2010 (PDT)

A BIG power bump this morning at about 02:45 knocked everything—RF, main magnet, vacuum—off-line. ISAC had a slew of failures as well. Recovery is underway, but we don't know how long it's going to take. Check the cyclotron and ISAC logs for more information:

https://cyc-elog.triumf.ca/

https://isac-elog.triumf.ca/

We'll update things as we know more. — Morton

[edit] 20:44, 7 April 2010 (PDT)

Last night's ISIS development ran much longer than expected and resulted in a need for a fair bit of cyclotron tuning. That's why protons weren't restored until this morning. Now that they have been, we're delivering 105/70/55 µA to BL1A/2A/2C4, respectively with µSR experiments running on M9 and M20. PiENu continues to run on M13 while M15 is still being used for testing.

8Li is in the process of being delivered to the βNQR facility at ISAC (about eight hours later than we'd hoped, but still—barely—according to the published schedule). That experiment will run until Monday morning. Otherwise no changes to RIB delivery are expected. The DRAGON group will take stable beam (16O) from OLIS over the weekend to investigate the acceptance of their spectrometer. — Morton

[edit] 20:58, 6 April 2010 (PDT)

The first long maintenance/ISIS filament change of the schedule is drawing to a close. The plan was to implement a new theoretical tune for ISIS before injection and beam delivery later this evening. Some proton beam is expected overnight.

ISAC is between experiments. Tonight's plan is to restore and check the tune to the yield station once protons are restored so that the yields of Li isotopes can be checked tomorrow morning. Once that's complete, the tune to βNQR will be restored and checked before delivery to the next experiment (hopefully by mid-day or so). — Morton

[edit] 07:43, 25 March 2010 (PDT)

All three beamlines are now online: BL1A (~105 µA), 2A (10 µA), and 2C4 (~35 µA). The beam energy in both 1A and 2A is 480 MeV. The 2A and 2C4 currents will both be increased over the next few days as targets improve, etc.

A target scan on BL2A will likely be carried out at some point today. Otherwise beam tuning is still in progress. — Morton

[edit] 20:43, 23 March 2010 (PDT)

Beam is currently running down BL1A and 2C4 (~100 µA and 25 µA, respectively). The 2C4 current is limited by STF water conductivity and will be increased as that improves. ISAC is not expected to request protons down BL2A until late Wednesday at the earliest.

The cyclotron is currently delivering protons at 500 MeV. Depending on how startup proceeds, the beam energy may be reduced to 480 MeV for the start of this running period in order to test the machine stability and production yields on the different beamlines at this energy. Running at 480 MeV reduces the cyclotron spills (as the H- beam makes fewer turns at high energy) which, in turn, reduces activation due to scattered beam. If we can run at 480 MeV on a routine basis, it would reduce the fields in the vault during shutdown allowing for work to begin sooner in shutdown with less dose to those doing the work. — Morton

[edit] 16:17, 18 March 2010 (PDT)

And we're back.

We're approaching the end of shutdown. The cyclotron is under vacuum, the main magnet is on, and both the RF and the inflector are at or near operating voltage. ISIS is running and the tuning effort is underway; the hope is to inject beam into the cyclotron within the next couple of days. Extraction down BL1A will follow, with extraction down BL2A planned for next week.

Regular beam delivery reports will commence next week. — Morton

Personal tools